# An Efficient Reservation Protocol for Medium Access: When Tree Splitting Meets Reinforcement Learning

### v1

@@ I Introduction
Access algorithms primarily operate at the medium access control (MAC) layer. Early approaches include ALOHA [9], where terminals retransmit after a random delay if a collision occurs. A variant called slotted ALOHA [10] was later proposed, where time is globally synchronized and divided into slots, and transmission occurs only at the beginning of a time slot. Another important class of access algorithms is the tree algorithm [11, 12], where terminals are assigned hierarchical levels based on past transmission outcomes, with subsequent decisions made according to the assigned level. Systematic overviews of ALOHA-like and tree-based algorithms can be found in [13, 14]. The performance of the above algorithms and the theoretical limits of such multiple access algorithms have been widely studied [15, 16, 17, 18, 19, 20]. However, the direct data transmission mechanism in the above algorithms can lead to excessive resource waste during collisions, especially when data packets are large. Therefore, researchers have explored alternatives such as carrier sense multiple access (CSMA) [21, 22], where terminals can detect whether the channel is busy for a short period before transmitting. Leveraging this information, terminals wait when the channel is busy to avoid collisions and initiate transmission as soon as the channel becomes idle. However, collisions can still occur when multiple terminals detect the idle channel and initiate transmission simultaneously. To further reduce collisions during data transmission, reservation-based protocols [8] have been developed. These protocols use short reservation packets to reserve the communication channel before data transmission. This process, known as the contention cycle, improves performance by separating contention resolution from data transmission. For time-sensitive applications where data must be transmitted within a strict time frame, polling algorithms [23] provide a structured approach by having a central terminal periodically poll devices to determine transmission needs. Beyond collision avoidance, random multiple-access (RMA) has also been investigated through interdisciplinary perspectives.
@@
One prominent approach is information-theoretic RMA that integrates information theory, which traditionally ignores random packet arrivals, and networking, which traditionally lacks detailed physical-layer modeling [24, 25]. For instance, the authors in [26] leverage the adaptive coding schemes to improve throughput. The random access problem is also formulated as a coding problem in [27]. Under this formulation, [28] introduces coded slotted ALOHA (CAS), which combines packet erasure correcting codes and successive interference cancellation (SIC). A systematic review of coding-theoretic RMA approaches can be found in [29]. Another key direction applies game theory to model the RMA problem as a non-cooperative game, where terminals make independent access decisions based on their strategies and payoffs. This approach balances fairness and efficiency in channel access [30, 31, 32]. The RMA problem has also been examined as a network utility maximization (NUM) problem [33, 34, 35], emphasizing global fairness. While these approaches have advanced RMA optimization, communication complexity [36] remains high, particularly under heavy traffic conditions. To address this challenge, re-examining RMA through emerging technologies, such as artificial intelligence (AI), presents a promising direction for future research.
@@
  • This paper investigates the RMA problem in a setting where a reservation protocol governs channel reservation before data transmission. Drawing inspiration from task-oriented communication and leveraging the structural properties of the problem, we formulate the RMA problem as a POMDP. Based on this formulation, we propose a reservation protocol that optimizes the reservation process through carefully designed interaction mechanisms for information exchange.
  • We consider a general class of reservation protocols where terminals with identical information adopt identical probabilistic actions. Within this class, we propose an efficient reservation protocol that is learned through real-time dynamic programming (RTDP). Additionally, we introduce a pre-training strategy for the algorithm that significantly accelerates the learning process. The proposed reservation protocol also ensures first-in-first-out (FIFO) service, guaranteeing that packets arriving within the current contention cycle are always transmitted before those arriving in subsequent cycles. This structured transmission approach facilitates a more tractable analysis of system performance.
@@ II-A Reservation-Based Random Multiple Access
An illustration of the two strategies is provided in Fig. 1. For the remainder of this paper, we assume the distribution of $N_{k}$ is known prior to each channel reservation process.
@@ II-C Tree-Splitting-Based Reservation
According to [13], a reservation protocol is defined as a function of (1) the generation time of the data packet, (2) the channel feedback, and (3) the terminal’s decision history. Since no priority is assigned among terminals in this paper, the data packet generation time does not affect the reservation protocol. Moreover, as channel feedback is broadcast to all terminals without error, the reservation protocol relies solely on the terminal’s decision history. Therefore, we focus on the reservation protocols that exhibit the characteristics outlined above.
@@ V Simulation Results
where $k\leq N_{max}$. The initial belief state $b_{0}$, as required in Algorithm 1, is then given by $b_{0}(s)=P(\Gamma^{T}=k)$ where $s=G^{k}_{1}$. The operation of the system under the two frame division strategies is detailed as follows.
@@
In this case, channel reservation takes place at the start of each frame. If there is ongoing packet transmission at the beginning of the frame, it is paused to allow the channel reservation to proceed. Once the channel reservation concludes, packet transmission resumes. Note that all previously queued packets must be transmitted before the arrivals within the current frame are processed. This is achieved by having the terminal send a dedicated finish signal once it has finished transmitting its packet, and the corresponding channel feedback is broadcast to inform all terminals of the transmission status. By listening to the channel feedbacks, terminals can determine the number of pending transmissions in the queue and coordinate their transmission.
@@
In this case, a new frame begins only after all terminals with backlogged data have completed their transmissions. This is made possible by leveraging the finish signal described earlier. This strategy will be the default setting for the subsequent simulations.
@@
In the sequel, we study the performance of the proposed reservation protocol under various system settings, as visualized in Fig. 5. Fig. 5(a) shows that under light traffic conditions, all frame division strategies achieve similar performance. However, as traffic intensity increases, performance degrades when the frame length is fixed. For short frames, the degradation is primarily due to the frequent reservation process, which consumes excessive channel resources. In the case of long frames, the performance degradation primarily results from inefficient channel utilization, as the number of idle slots at the end of the frame increases. The dynamic frame division strategy, on the other hand, performs better under heavy traffic, as reservations are initiated immediately after data transmissions complete. The average delay under different frame division strategies is shown in Fig. 5(b). As we can see, for fixed frames, the delay dramatically increases when the arrival rate exceeds a certain threshold. This occurs because overly frequent reservations in a fixed frame lead to severe data backlog. In contrast, for the dynamic frame, the delay remains lower and stable, as data transmission proceeds uninterrupted by the reservation process. It is also worth noting that the proposed reservation protocol ensures FIFO service. This is in contrast to the benchmark protocols shown in Fig. 4, where the delivery order is not guaranteed. The effective throughput under different data-to-overhead ratios is shown in Fig. 5(c). As shown, the proposed reservation protocol achieves effective throughput very close to the upper limit, which is indicated by the dotted lines. Note that some data points exceed the upper limit due to the stochastic data arrival. The average delay when the data-to-overhead ratio varies is shown in Fig. 5(d). As expected, the delay increases with the data-to-overhead ratio. This is because larger data packets require longer transmission times. Additionally, for a fixed data-to-overhead ratio, an increase in the arrival rate leads to longer waiting times due to the larger number of packet arrivals and the FIFO service provided by the proposed reservation protocol.
@@
Finally, we investigate the impact of RTDP-Bel’s parameters on its performance. The results are shown in Fig. 6. In these simulations, we run Algorithm 1 for multiple trials and omit stochastic data arrival. Hence, we omit the data transmission process and fix the initial belief state as $b_{0}=[0.1,0.1,0.3,0.3,0.2]$. The first parameter we investigate is the quantization parameter $q$. From Fig. 6(a), we can see that a small $q$ leads to higher cost due to the merging of distinct states with different optimal actions. However, increasing $q$ from $1$ to $20$ enlarges the size of the hash table from $1194$ to $3751$ entries, but the performance improvement is marginal. A similar trend is also observed in Fig. 6(b), which presents the long-term behavior. Next, we examine the effect of the discretization parameter $d$. The short-term and long-term behavior are shown in Fig. 6(c) and Fig. 6(d), respectively. We observe that a small $d$ leads to performance degradation, while increasing $d$ improves performance. However, the performance gain is relatively small compared to the significant increase in computational complexity. For instance, increasing $d$ from $5$ to $15$ results in an eightfold increase in the size of the decision variable, which considerably raises the computational demand. Fig. 6 also highlights the benefits of pre-training. Specifically, the ”No Pre-training” case exhibits high costs, whereas the proposed reservation protocol with pre-training exhibits reduced costs. Additionally, in long-term behavior, the proposed reservation protocol with $d=15$ achieves an average cost of around $7.1$, compared to $5.4$ for the optimal genie-aided policy. This performance gap arises from the lack of knowledge about the number of active terminals. Hence, the extra cost is required to confirm that there are no active terminals. Since this extra cost is around $1.7$, and confirmation incurs a cost of at least $1$, the proposed reservation protocol performs close to the optimal genie-aided policy. However, the genie-aided performance relies on the knowledge about the active terminals’ distribution.

### v2

@@ I Introduction
Access algorithms primarily operate at the medium access control (MAC) layer. Early approaches include ALOHA [9], where terminals retransmit after a random delay if a collision occurs. A variant called slotted ALOHA [10] was later proposed, where time is globally synchronized and divided into slots, and transmission occurs only at the beginning of a time slot. Another important class of access algorithms is the tree algorithm [11, 12], where terminals are assigned hierarchical levels based on past transmission outcomes, with subsequent decisions made according to the assigned level. Systematic overviews of ALOHA-like and tree-based algorithms can be found in [13, 14]. The performance of the above algorithms and the theoretical limits of such multiple access algorithms have been widely studied [15, 16, 17, 18, 19, 20]. However, the direct data transmission mechanism can lead to excessive resource waste during collisions, especially when data packets are large. Therefore, researchers have explored alternatives such as carrier sense multiple access (CSMA) [21, 22], where terminals can detect whether the channel is busy for a short period before transmitting. Leveraging this information, terminals wait when the channel is busy to avoid collisions and initiate transmission as soon as the channel becomes idle. However, collisions can still occur when multiple terminals detect the idle channel simultaneously. To further reduce collisions during data transmission, reservation-based protocols [8] have been developed. These protocols use short reservation packets to reserve the communication channel before data transmission. This process, known as the contention cycle, improves performance by separating contention resolution from data transmission. For time-sensitive applications where data must be transmitted within a strict time frame, polling algorithms [23] provide a structured approach by having a central terminal periodically poll devices to determine transmission needs. Beyond collision avoidance, random multiple-access (RMA) has also been investigated through interdisciplinary perspectives.
@@
One prominent approach is information-theoretic RMA that integrates information theory, which traditionally ignores random packet arrivals, and networking, which traditionally lacks detailed physical-layer modeling [24, 25]. For instance, the authors in [26] leverage the adaptive coding schemes to improve throughput. The random access problem is also formulated as a coding problem in [27]. Under this formulation, [28] introduces coded slotted ALOHA (CAS), which combines packet erasure correcting codes and successive interference cancellation (SIC). A systematic review of coding-theoretic RMA approaches can be found in [29]. Another key direction applies game theory to model the RMA problem as a non-cooperative game, where terminals make independent access decisions based on their strategies and payoffs. This approach balances fairness and efficiency in channel access [30, 31, 32]. The RMA problem has also been examined as a network utility maximization (NUM) problem [33, 34, 35], emphasizing global fairness. While these approaches have advanced RMA optimization, communication complexity [36] remains high, particularly under heavy traffic conditions. Hence, re-examining RMA with emerging technologies, such as artificial intelligence (AI), shows great potential.
@@
  • This paper investigates the RMA problem in a setting where a reservation protocol governs channel reservation before data transmission. Leveraging the structural properties of the problem, we formulate the RMA problem as a POMDP. Based on this formulation, we propose a reservation protocol that optimizes the reservation process through carefully designed interaction mechanisms for information exchange. The reservation protocol is presented and optimized for the access stage in both wired and wireless local area networks (LANs). In particular, whether an ACK or NACK is feedback given the number of simultaneously attempting terminals, is the same as that assumed in wired LANs, such as Ethernet [43] and bus topology [44], as well as in ALOHA networks and the IEEE 802.11 DCF for wireless LANs (WLANs). As a result, the conceived protocol are compatible with the current hardware platform of IEEE 802.11, thereby being capable of replacing existing reservation schemes of IEEE 802.11 DCF without much hardware upgrades.
  • We consider a class of reservation protocols where terminals with identical information adopt identical probabilistic actions. Within this class, we propose an efficient reservation protocol that is learned through real-time dynamic programming (RTDP). Additionally, we introduce a pre-training strategy that significantly accelerates the learning. The proposed reservation protocol also ensures first-in-first-out (FIFO) service, guaranteeing that packets are transmitted in the order of their arrival contention cycles. This structured transmission facilitates a more tractable performance analysis.
@@ II-A Reservation-Based Random Multiple Access
An illustration of the two strategies is provided in Fig. 1. For the remainder of this paper, we assume the distribution of $N_{k}$ is known prior to each channel reservation process. It can be noted that we are interested in the access stage of the reservation-based multiaccess communications. In wireless environments, the physical-layer settings can be treated in the same manner as those in the IEEE 802.11 DCF, thereby allowing related issues to be handled similarly. In this context, the reservation process can be modeled from an information-theoretic perspective, as detailed in Section II-B.
@@ II-C Tree-Splitting-Based Reservation
According to [13], a reservation protocol is defined as a function of (1) the generation time of the data packet, (2) the channel feedback, and (3) the terminal’s decision history. Since no priority is assigned among terminals in this paper, the data packet generation time does not affect the reservation protocol. Moreover, as channel feedback is broadcast to all terminals without error, the reservation protocol relies solely on the terminal’s decision history. Therefore, we focus on the reservation protocol outlined above.
@@ V Simulation Results
where $k\leq N_{max}$. The initial belief state $b_{0}$, as required in Algorithm 1, is then given by $b_{0}(s)=P(\Gamma^{T}=k)$ where $s=G^{k}_{1}$. Throughout our simulations, the ACK/NACK feedback, given the number of simultaneously attempting terminals, is the same as that assumed in the IEEE 802.11 DCF and ALOHA. More specifically, an ACK is broadcast only when a single terminal attempts transmission, and a NACK is broadcast if more than one terminal attempts transmission. Then, the operation of the system under the two frame division strategies is detailed as follows.
@@
Channel reservation takes place at the start of each frame. If there is ongoing packet transmission at the beginning of the frame, it is paused to allow the channel reservation to proceed. Once the channel reservation concludes, packet transmission resumes. Note that all previously queued packets must be transmitted before the arrivals within the current frame. This is achieved by having the terminal send a finish signal once it has finished transmitting, and the corresponding channel feedback is broadcast to inform all terminals of the transmission status. By listening to the channel feedbacks, terminals can determine the number of pending transmissions in the queue and coordinate their transmission.
@@
A new frame begins only after all terminals with backlogged data have completed their transmissions. This is made possible by leveraging the finish signal described earlier. This strategy will be the default setting for the subsequent simulations.
@@
In the sequel, we study the performance of the proposed reservation protocol under various system settings, as visualized in Fig. 5. Fig. 5(a) shows that under light traffic conditions, all frame division strategies achieve similar performance. However, as traffic intensity increases, performance degrades when the frame length is fixed. For short frames, the degradation is primarily due to the frequent reservation process, which consumes excessive channel resources. For long frames, the performance degradation primarily results from inefficient channel utilization, as the number of idle slots increases. The dynamic frame division strategy, on the other hand, performs better under heavy traffic, as reservations are initiated immediately after data transmissions complete. The average delay under different frame division strategies is shown in Fig. 5(b). As we can see, for fixed frames, the delay dramatically increases when the arrival rate exceeds a certain threshold. This occurs because overly frequent reservations in a fixed frame lead to severe data backlog. In contrast, for the dynamic frame, the delay remains lower and stable, as data transmission proceeds uninterrupted by the reservation process. It is also worth noting that the proposed reservation protocol ensures FIFO service. This is in contrast to the benchmark protocols considered in Fig. 4, where the delivery order is not guaranteed. The effective throughput under different $\rho$ is shown in Fig. 5(c). As shown, the proposed reservation protocol achieves effective throughput very close to the upper limit. Note that some data points exceed the upper limit due to the stochastic data arrival. The average delay when $\rho$ varies is shown in Fig. 5(d). As expected, the delay increases with $\rho$. This is because larger data packets require longer transmission times. Additionally, for a fixed $\rho$, an increase in the arrival rate leads to longer waiting times due to the FIFO service provided by the proposed reservation protocol.
@@
Finally, we investigate the impact of RTDP-Bel’s parameters on its performance. The results are shown in Fig. 6. In these simulations, we run Algorithm 1 for multiple trials and omit stochastic data arrival. Hence, we omit the data transmission process and fix the initial belief state as $b_{0}=[0.1,0.1,0.3,0.3,0.2]$. The first parameter we investigate is the quantization parameter $q$. From Fig. 6(a), we can see that a small $q$ leads to higher cost due to the merging of distinct states with different optimal actions. However, increasing $q$ from $1$ to $20$ enlarges the size of the hash table from $1194$ to $3751$ entries, but the performance improvement is marginal. A similar trend is also observed in Fig. 6(b), which presents the long-term behavior. Next, we examine the effect of the discretization parameter $d$. The short-term and long-term behavior are shown in Fig. 6(c) and Fig. 6(d), respectively. We observe that a small $d$ leads to performance degradation, while increasing $d$ improves performance. However, the performance gain is relatively small compared to the significant increase in computational complexity. For instance, increasing $d$ from $5$ to $15$ results in an eightfold increase in the size of the decision variable. Fig. 6 also highlights the benefits of pre-training. Specifically, the ”No Pre-training” case exhibits high costs, whereas the proposed reservation protocol with pre-training exhibits reduced costs. Additionally, in long-term behavior, the proposed reservation protocol with $d=15$ achieves an average cost of around $7.1$, compared to $5.4$ for the optimal genie-aided policy. This performance gap arises from the lack of knowledge about the number of active terminals. Hence, the extra cost is required to confirm that there are no active terminals. Since this extra cost is around $1.7$, and confirmation incurs a cost of at least $1$, the proposed reservation protocol performs close to the optimal genie-aided policy.